# Bulk viscosity of nuclear matter with pions in the neutrino-trapped regime

### v1

@@ Abstract
Recent work [B. Fore and S. Reddy, Phys. Rev. C 101 035809 (2020)] has shown that the population of thermal pions could modify the equation of state and transport properties of hot and dense neutron-rich matter and introduce new reaction pathways to change the proton fraction. In this article we study their impact on the bulk viscosity of dense matter, focusing on the neutrino-trapped regime that would be realized in neutron star mergers and supernovae. We find that the presence of a thermal population of pions alters the bulk viscosity by modifying the EoS (via the susceptibilities) and by providing new reaction pathways to achieve beta-equilibrium. In neutron star merger conditions, the bulk viscosity in neutrino-trapped $npe\mu$ matter (without pions) has its peak at temperatures of at most a couple MeV and is quite small at temperatures of tens of MeV. We find that thermal pions enhance the low-temperature peak of the bulk viscosity by a factor of a few and shift it to slightly lower temperatures. At higher temperatures, where the pion abundance is large but the bulk viscosity is traditionally small, pions can increase the bulk viscosity by an order of magnitude or more, although it is still orders of magnitude smaller than its peak value.
@@ I Introduction
In this work, we study the effect of thermal pions on bulk viscosity, considering their effects on out-of-equilibrium chemical reactions and the EoS. We will fully develop the formalism for bulk viscosity in neutrino-trapped nuclear matter containing neutrons, protons, electrons, muons, and a thermal population of pions. In $npe$ matter, the proton fraction is adequate to specify the composition since the neutron and electron densities can be determined via baryon number conservation and charge neutrality. There is one chemical equilibration channel and therefore the bulk viscosity exhibits a single resonant peak. For a given density, the resonance occurs at a temperature when the equilibration rate becomes comparable to the oscillation frequency [28, 9]. Systems with several independent particle species can have a bulk viscosity with a complicated temperature dependence with several overlapping or distinct resonances. The rule of thumb is one resonance for each independent equilibration channel. This complexity has been studied in earlier work for the case of $npe\mu$ matter [29, 4, 8, 14], and in quark matter [30, 31, 32, 33, 34, 13]. Here we perform a similar analysis for $npe\mu$ matter containing a population of thermal pions, taking into account the effect of pions on the out-of-equilibrium reactions and the thermodynamic properties such as the susceptibilities that influence the bulk viscosity.
@@
The population of negatively charged thermal pions in neutron-rich matter at high density is enhanced due to a rapid increase in their associated chemical potential. The negative charge chemical potential, denoted by $\hat{\mu}=\mu_{n}-\mu_{p}$ where $\mu_{n}$ and $\mu_{p}$ are the neutron and proton chemical potentials, respectively, is needed to ensure electric charge neutrality and its magnitude (which depends on the nuclear symmetry energy) becomes comparable to the vacuum pion mass $m_{\pi}$. Consequently, in the absence of interactions, the pion number density increases exponentially as $\exp{(\hat{\mu}-m_{\pi})/T}$. Calculations in [27], which includes pion-nucleon interactions that are based on the virial expansion, found that interactions increase the pion fraction at finite temperatures. This study found that although s-wave interactions are repulsive, a strong and attractive p-wave interaction can greatly enhance the pion number at a modest temperature, motivating the study of the presence of pions on properties of hot, dense matter.
@@
To model interactions between pions and nucleons we use a construction of the pion self-energy, $\Sigma_{\pi^{-}}(p)$ which is structurally similar to the one introduced in [27]. The pion dispersion relation is given by
@@
$V^{(ps)}_{N\pi^{-}}(p_{\rm cm})=-\sum_{I,l,\nu}\alpha_{l}(2l+1)\frac{2\pi\delta_{l,\nu}^{I}}{{\bar{m}}~{}p_{\rm cm}}$
@@
This modified pseudopotential now has two parameters $\alpha_{l}$ rather than the single value used in [27]. This change is motivated by the fact that the model under-represents the s-wave repulsion present in pion-nucleon interactions by suppressing the s-wave and p-wave interactions equally. To prevent this we set the parameter $\alpha_{0}$=1 and fit the value of the remaining parameter $\alpha_{1}$ such that the model produces the same pion fraction as the virial calculation presented in [27] at a baryon density equal to nuclear saturation density, a temperature of 30 MeV, and with both electron-type and muon-type conserved lepton fractions equal to 0.1. Unlike in [27], where the value of $\alpha$ was explored at various temperatures and densities, our fitting parameter, $\alpha_{1}$, has a fixed value of 0.23596 for all conditions considered.
@@
Figure 1: The pion dispersion relation in vacuum (dashed line) and in dense matter, as predicted by our EoS (solid lines). The pion “mass” [$E(p=0)$] is shifted from its vacuum value (139.6 MeV) to 177.4 MeV at $1n_{0}$ and to 194.0 MeV at $1.6n_{0}$. The temperature is taken to be 1 MeV.
@@
The pion dispersion relation predicted by Eq. 1 is shown in Fig. 1. As noted earlier, the shape of the dispersion curve can be understood as arising due to the competition between s-wave repulsion and strong p-wave attraction. The self-energy at zero momentum leads to a positive shift of the pion mass of about $40$ MeV at $n_{B}=n_{0}$ and by about $54$ MeV at $n_{B}=1.6~{}n_{0}$. These large shifts are consistent with recent calculations of the pion mass in dense neutron-rich matter reported in Ref. [36]. With increasing momentum, the self-energy becomes negative due to attractive p-wave interactions and produces the distinct non-monotonic behavior characterized by the maximum and minimum seen in the figure. Since the typical pion momentum $p_{\pi}\simeq 3T$, with increasing temperature the low lying states at higher momentum will be preferentially occupied and the non-monotonic dispersion relation leads to a stronger temperature dependence of the thermal pion population as will be discussed below.
@@
Since we aim to cover the neutrino-trapped conditions that exist in supernova and neutron star mergers, we choose a variety of $Y_{L,i}$ to mimic these conditions. No neutron star merger simulation has, to this point, dynamically included muons, and therefore we refer to a post-processing analysis [37] which predicts that matter above saturation density could have $Y_{Le}$ in the range of 0.04-0.10 and $Y_{L\mu}$ in the range 0.01 - 0.07. Supernovae simulations reliably predict $Y_{Le}$ to be between 0.3 and 0.4 [38]. Muons have been included in recent supernovae simulations [39, 40], and since the net muon number $Y_{L\mu}$ in the supernova progenitor is close to zero, it remains that way throughout the supernova. To study these different physical situations, we will focus on the two configurations $\{Y_{Le},Y_{L\mu}\}=\{0.05,0\}$ (a good description of neutron star merger conditions) and $\{Y_{Le},Y_{L\mu}\}=\{0.3,0\}$ (a good approximation of supernova conditions). In the merger case, the two merging neutron stars are cold and do not contain a large muon population. Upon rapid heating at the contact of the two stars, the muon number is frozen due to the trapping of neutrinos. For supernovae, the matter is initially at low density and therefore does not contain a large muon population, and once the matter becomes dense enough to trap neutrinos, the muon number becomes frozen in. While it is almost certainly not zero in either physical situation, it is reasonable to choose zero as a representative case. In the main text, we discuss the particle content, susceptibilities, equilibration rates, and bulk viscosity for matter with the above two choices of lepton fractions, and put the corresponding results for equal values of the conserved electron and muon fraction $\{Y_{Le},Y_{L\mu}\}=\{0.05,0.05\}\text{ and }\{0.3,0.3\}$ in Appendix A.
@@ III Bulk viscosity
where $\delta n_{B}\ll n_{B}$ and we have chosen $\delta n_{B}$ to be real. A change in density can push the nuclear matter fluid element out of beta equilibrium by an amount $\delta\mu$ (to be precisely defined later in Eqs. 14a-14c). In this study, we assume the density oscillation amplitude is small enough to only push the system slightly out of chemical equilibrium, such that $\delta\mu\ll T$. The resulting bulk viscosity calculated under these conditions is called “subthermal” [43]. Below, we extend the subthermal bulk viscosity derivation in $npe$ matter with one equilibrating chemical potential $\delta\mu$ [28, 44, 14] to neutrino-trapped $npe\mu\pi^{-}$ matter, which will turn out to have three equilibrating chemical potentials $\left\{\delta\mu_{1},\delta\mu_{2},\delta\mu_{3}\right\}$.
@@
The $\Im{(\delta P)}$ term that represents the phase lag between the pressure $P(t)$ and the baryon density $n_{B}(t)$ which will give rise to bulk viscous energy dissipation via the mechanism of $p\mathop{dV}$ work.
@@
$\dfrac{\mathop{dn_{B}}}{\mathop{dt}}+n_{B}\nabla\cdot\mathbf{v}=0,$ (6)
@@
$\left\langle\dfrac{\mathop{d\varepsilon}_{\text{osc}}}{\mathop{dt}}\right\rangle=-\dfrac{1}{2}\left(\dfrac{\delta n_{B}}{n_{B}}\right)^{2}\omega^{2}\zeta.$ (7)
The energy lost due to $p\mathop{dV}$ work, averaged over one cycle is
$\left\langle\dfrac{\mathop{d\varepsilon}_{\text{diss}}}{\mathop{dt}}\right\rangle$ $=\dfrac{\omega}{2\pi}\int_{0}^{2\pi/\omega}\mathop{dt}\dfrac{P(t)}{n_{B}(t)}\dfrac{\mathop{dn_{B}}(t)}{\mathop{dt}}$ (8)
$=\dfrac{\omega^{2}}{2\pi}\left(\dfrac{\delta n_{B}}{n_{B}}\right)\Im{(\delta P)}\int_{0}^{2\pi/\omega}\mathop{dt}\sin^{2}{\left(\omega t\right)}$
$=\dfrac{\omega}{2}\left(\dfrac{\delta n_{B}}{n_{B}}\right)\Im{(\delta P)},$
where we used the Eq. 4 for the time-dependence of the pressure. The energy lost by the oscillation is gained by the fluid element, and thus Eq. 7 and 8 sum to zero and we find
@@
$P=P_{0}+\dfrac{\partial P}{\partial n_{B}}\bigg{|}_{T,x_{p},x_{\mu},x_{\pi}}\Re{(\delta n_{B}e^{i\omega t})}$ (11)
$+\dfrac{\partial P}{\partial x_{p}}\bigg{|}_{T,n_{B},x_{\mu},x_{\pi}}\Re{(\delta x_{p}e^{i\omega t})}+\dfrac{\partial P}{\partial x_{\mu}}\bigg{|}_{T,n_{B},x_{p},x_{\pi}}\Re{(\delta x_{\mu}e^{i\omega t})}$
$+\dfrac{\partial P}{\partial x_{\pi}}\bigg{|}_{T,n_{B},x_{p},x_{\mu}}\Re{(\delta x_{\pi}e^{i\omega t})}.$
Evidently (from Eqs. 4, 10, and 11)
$\Im{(\delta P)}$ $=\dfrac{\partial P}{\partial x_{p}}\bigg{|}_{T,n_{B},x_{\mu},x_{\pi}}\Im{(\delta x_{p})}$ (12)
$+\dfrac{\partial P}{\partial x_{\mu}}\bigg{|}_{T,n_{B},x_{p},x_{\pi}}\Im{(\delta x_{\mu})}+\dfrac{\partial P}{\partial x_{\pi}}\bigg{|}_{T,n_{B},x_{p},x_{\mu}}\Im{(\delta x_{\pi})},$
@@ III.1 Classification of chemical reactions in neutrino-trapped $npe\mu\pi^{-}$ matter
There are six classes [^1] Reactions within each class are related by “crossing” (e.g. direct Urca neutron decay and electron capture) or by the addition of a spectator particle to both sides of the reaction (for example, direct Urca neutron decay and modified Urca neutron decay). of weak reactions that can occur in neutrino-trapped $npe\mu\pi$ matter, and therefore six chemical potential differences $\delta\mu_{i}$ that are zero in chemical equilibrium. However, three of them are redundant and can be written in terms of three independent equilibrating chemical potentials
@@
We assume in this analysis that particles are chemically equilibrated with their respective antiparticles, $\mu_{X}=-\mu_{\bar{X}}$. This is a good assumption for the hadrons and charged leptons, though for neutrinos it is merely a simplifying assumption [^2] If neutrinos and antineutrinos were not equilibrated with each other, as is likely the case except at quite large temperatures, we would have to assign distinct chemical potentials $\mu_{\nu_{e}},\mu_{\bar{\nu}_{e}},\mu_{\nu_{\mu}},\mu_{\bar{\nu}_{\mu}}$ and consider reactions such as $n+n\leftrightarrow n+n+\nu_{e}+\bar{\nu}_{e}$ or $\nu_{e}+\bar{\nu}_{e}\leftrightarrow\nu_{\mu}+\bar{\nu}_{\mu}$.. Given this, the six classes of reactions are listed in Table 1 below.
@@
6) $\delta\mu_{6}=\mu_{\mu}+\mu_{\nu_{e}}-\mu_{e}-\mu_{\nu_{\mu}}=\delta\mu_{1}-\delta\mu_{2}$ $\mu^{-}\leftrightarrow e^{-}+\bar{\nu_{e}}+\nu_{\mu}$ $\lambda_{i}(\delta\mu_{1}-\delta\mu_{2})$
@@
The process $n\leftrightarrow p+\pi^{-}$ needs a spectator to proceed, but even the spectator reaction $n+n\leftrightarrow n+p+\pi^{-}$ is very fast compared to all other timescales in the problem as long as $T\gtrsim 1\text{ MeV}$, so in our numerical calculations we will take the $\lambda_{3}\rightarrow\infty$ limit and never have to calculate the rate of $n+n\leftrightarrow n+p+\pi^{-}$. The rate of this process can be calculated using models such as the one developed in [56].

### v3

@@ Abstract
Recent work [B. Fore and S. Reddy, Phys. Rev. C 101 035809 (2020)] has shown that a population of thermal pions could modify the equation of state and transport properties of hot and dense neutron-rich matter and introduce new reaction pathways to change the proton fraction. In this article we study their impact on the bulk viscosity of dense matter, focusing on the neutrino-trapped regime that would be realized in neutron star mergers and supernovae. We find that the presence of a thermal population of pions alters the bulk viscosity by modifying the EoS (via the susceptibilities) and by providing new reaction pathways to achieve beta equilibrium. In neutron star merger conditions, the bulk viscosity in neutrino-trapped $npe\mu$ matter (without pions) has its peak at temperatures of at most a couple MeV and is quite small at temperatures of tens of MeV. We find that thermal pions enhance the low-temperature peak of the bulk viscosity by a factor of a few and shift it to slightly lower temperatures. At higher temperatures, where the pion abundance is large but the bulk viscosity is traditionally small, pions can increase the bulk viscosity by an order of magnitude or more, although it is still orders of magnitude smaller than its peak value.
@@ I Introduction
In this work, we study the effect of thermal pions on bulk viscosity, considering their effects on out-of-equilibrium chemical reactions and on the EoS. We will fully develop the formalism for bulk viscosity in neutrino-trapped nuclear matter containing neutrons, protons, electrons, muons, and a thermal population of pions. In $npe$ matter, the proton fraction is adequate to specify the composition since the neutron and electron densities can be determined via baryon number conservation and charge neutrality. There is one chemical equilibration channel and therefore the bulk viscosity exhibits a single resonant peak. For a given density, the resonance occurs at a temperature when the equilibration rate becomes comparable to the oscillation frequency [28, 9]. Systems with several independent particle species can have a bulk viscosity with a complicated temperature dependence with several distinct or overlapping resonances. The rule of thumb is one resonance for each independent equilibration channel. This complexity has been studied in earlier work for the case of $npe\mu$ matter [29, 4, 8, 14], and in quark matter [30, 31, 32, 33, 34, 13]. Here we perform a similar analysis for $npe\mu$ matter containing a population of thermal pions, taking into account the effect of pions on the out-of-equilibrium reactions and the thermodynamic properties such as the susceptibilities that influence the bulk viscosity.
@@
The population of negatively charged thermal pions in neutron-rich matter at high density is enhanced due to a rapid increase in their associated chemical potential. The negative charge chemical potential, denoted by $\hat{\mu}=\mu_{n}-\mu_{p}$ where $\mu_{n}$ and $\mu_{p}$ are the neutron and proton chemical potentials, respectively, is needed to ensure electric charge neutrality and its magnitude (which depends on the nuclear symmetry energy) becomes comparable to the vacuum pion mass $m_{\pi}$. Consequently, in the absence of interactions, the pion number density increases exponentially as $\exp{\left[(\hat{\mu}-m_{\pi})/T\right]}$. Calculations in [27], which includes pion-nucleon interactions that are based on the virial expansion, found that interactions increase the pion fraction at finite temperatures. This study found that although s-wave interactions are repulsive, a strong and attractive p-wave interaction can greatly enhance the pion number at a modest temperature, motivating the study of the presence of pions on properties of hot, dense matter.
@@
The neutron and proton sector of the equation of state is described here with a nonrelativistic Skyrme model called NRAPR, described in [27, 36]. To model interactions between pions and nucleons we use a construction of the pion self-energy, $\Sigma_{\pi^{-}}(p)$ which is structurally similar to the one introduced in [27]. The pion dispersion relation is given by
@@
$V^{(ps)}_{N\pi^{-}}(p_{\rm cm})=-\sum_{I,l,\nu}\alpha_{l}(2l+1)\frac{2\pi\delta_{l,\nu}^{I}}{{\bar{m}}~{}p_{\rm cm}},$
@@
where the sum is over the isospin $I$, the angular momentum $l$, and the nucleon spin $\nu$. The symbols $\bar{m}$ and $p_{\text{cm}}$ represent the reduced mass of the pion-nucleon system and the center of mass momentum, respectively. The term $\delta_{l,\nu}^{I}$ is the phase shift in the given channel. The modified pseudopotential now has two parameters $\alpha_{l}$ rather than the single value used in [27]. This change is motivated by the fact that the model under-represents the s-wave repulsion present in pion-nucleon interactions by suppressing the s-wave and p-wave interactions equally. To prevent this we set the parameter $\alpha_{0}$=1 and fit the value of the remaining parameter $\alpha_{1}$ such that the model produces the same pion fraction as the virial calculation presented in [27] at a baryon density equal to nuclear saturation density, a temperature of 30 MeV, and with both electron-type and muon-type conserved lepton fractions equal to 0.1. Unlike in [27], where the value of $\alpha$ was explored at various temperatures and densities, our fitting parameter, $\alpha_{1}$, has a fixed value of 0.23596 for all conditions considered.
@@
Figure 1: The pion dispersion relation in vacuum (dashed line) and in dense matter, as predicted by our EoS (solid lines). The pion “mass” [$E_{\pi^{-}}(p=0)$] is shifted from its vacuum value (139.6 MeV) to 177.4 MeV at $1n_{0}$ and to 194.0 MeV at $1.6n_{0}$. The temperature is taken to be 1 MeV.
@@
The pion dispersion relation predicted by Eq. 1 is shown in Fig. 1. As noted earlier, the shape of the dispersion curve can be understood as arising due to the competition between s-wave repulsion and strong p-wave attraction. The self-energy at zero momentum leads to a positive shift of the pion “mass” [$E_{\pi^{-}}(p=0)$] of about $40$ MeV at $n_{B}=n_{0}$ and by about $54$ MeV at $n_{B}=1.6~{}n_{0}$. These large shifts are consistent with recent calculations of the pion mass in dense neutron-rich matter reported in Ref. [37]. With increasing momentum, the self-energy becomes negative due to attractive p-wave interactions and produces the distinct non-monotonic behavior characterized by the maximum and minimum seen in the figure. Since the typical pion momentum $p_{\pi}\simeq 3T$, with increasing temperature the low lying states at higher momentum will be preferentially occupied and the non-monotonic dispersion relation leads to a stronger temperature dependence of the thermal pion population as will be discussed below.
@@
Since we aim to cover the neutrino-trapped conditions that exist in supernovae and neutron star mergers, we choose a variety of $Y_{L,i}$ to mimic these conditions. No neutron star merger simulation has, to this point, dynamically included muons, and therefore we refer to a post-processing analysis [38] which predicts that matter above saturation density could have $Y_{Le}$ in the range of 0.04-0.10 and $Y_{L\mu}$ in the range 0.01 - 0.07. Supernovae simulations reliably predict $Y_{Le}$ to be between 0.3 and 0.4 [39]. Muons have been included in recent supernovae simulations [40, 41], and since the net muon number $Y_{L\mu}$ in the supernova progenitor is close to zero, it remains that way throughout the duration of the supernova. To study these different physical situations, we will focus on the two configurations $\{Y_{Le},Y_{L\mu}\}=\{0.05,0\}$ (a good description of neutron star merger conditions) and $\{Y_{Le},Y_{L\mu}\}=\{0.3,0\}$ (a good approximation of supernova conditions). In the merger case, the two merging neutron stars are cold and do not contain a large muon population. Upon rapid heating at the contact of the two stars, the muon number is frozen due to the trapping of neutrinos. For supernovae, the matter is initially at low density and therefore does not contain a large muon population, and once the matter becomes dense enough to trap neutrinos, the muon number becomes frozen in. While it is almost certainly not zero in either physical situation, it is reasonable to choose zero as a representative case. In the main text, we discuss the particle content, susceptibilities, equilibration rates, and bulk viscosity for matter with the above two choices of lepton fractions, and put the corresponding results for equal values of the conserved electron and muon fraction $\{Y_{Le},Y_{L\mu}\}=\{0.05,0.05\}\text{ and }\{0.3,0.3\}$ in Appendix A.
@@ III Bulk viscosity
where $\delta n_{B}\ll n_{B}$ and we have chosen $\delta n_{B}$ to be real. Here, the real and imaginary parts of a quantity $X$ are denoted $\Re{(X)}$ and $\Im{(X)}$, respectively. A change in density can push the nuclear matter fluid element out of beta equilibrium by an amount $\delta\mu$ (to be precisely defined later in Eqs. 16a-16c). In this study, we assume the density oscillation amplitude is small enough to only push the system slightly out of chemical equilibrium, such that $\delta\mu\ll T$. The resulting bulk viscosity calculated under these conditions is called “subthermal” [44]. Below, we extend the subthermal bulk viscosity derivation in $npe$ matter with one equilibrating chemical potential $\delta\mu$ [28, 45, 14] to neutrino-trapped $npe\mu\pi^{-}$ matter, which will turn out to have three equilibrating chemical potentials $\left\{\delta\mu_{1},\delta\mu_{2},\delta\mu_{3}\right\}$.
@@
The $\Im{(\delta P)}$ term that represents the phase lag between the pressure $P(t)$ and the baryon density $n_{B}(t)$ which will give rise to bulk viscous energy dissipation via the mechanism of $PdV$ work.
@@
$\dfrac{\mathop{dn_{B}}(t)}{\mathop{dt}}+n_{B}(t)\nabla\cdot\mathbf{v}=0,$ (6)
@@
$\left\langle\dfrac{\mathop{d\varepsilon}_{\text{osc}}}{\mathop{dt}}\right\rangle$ $=-\dfrac{1}{2}\left(\dfrac{\delta n_{B}}{n_{B}(t)}\right)^{2}\omega^{2}\zeta$
$\approx-\dfrac{1}{2}\left(\dfrac{\delta n_{B}}{n_{B}}\right)^{2}\omega^{2}\zeta.$ (7)
The energy lost due to $PdV$ work is
$\mathop{dE_{\text{diss}}}=\mathop{d\left(\varepsilon V\right)}=-P\mathop{dV}.$ (8)
We can express this as the energy density lost due to $PdV$ work
$\mathop{d\varepsilon_{\text{diss}}}=\left(\dfrac{\varepsilon+P}{n_{B}}\right)\mathop{dn_{B}},$ (9)
where we used the fact that baryon number is conserved $\mathop{dN_{B}}=\mathop{d\left(n_{B}V\right)}=n_{B}\mathop{dV}+V\mathop{dn_{B}}=0$. Averaging over a cycle,
$\left\langle\dfrac{\mathop{d\varepsilon}_{\text{diss}}}{\mathop{dt}}\right\rangle$ $=\dfrac{\omega}{2\pi}\int_{0}^{2\pi/\omega}\mathop{dt}\left[\dfrac{P(t)+\varepsilon(t)}{n_{B}(t)}\right]\dfrac{\mathop{dn_{B}}(t)}{\mathop{dt}}$ (10)
$\approx\dfrac{\omega^{2}}{2\pi}\left(\dfrac{\delta n_{B}}{n_{B}}\right)\Im{(\delta P)}\int_{0}^{2\pi/\omega}\mathop{dt}\sin^{2}{\left(\omega t\right)}$
$\approx\dfrac{\omega}{2}\left(\dfrac{\delta n_{B}}{n_{B}}\right)\Im{(\delta P)},$
where we used Eq. 4 for the time-dependence of the pressure. We drop the energy density term, because the energy density oscillates in-phase with the baryon density and therefore drops out of the time-average [^1] This fact can be seen from the Taylor expansion of the pressure (Eq. 13). The partial derivatives of the pressure with respect to particle fractions have finite values (related to the susceptibilities, discussed later), which gives the pressure $P(t)$ terms that are proportional to sine which are thus of phase with the density oscillation. An analogous expansion of the energy density around beta equilibrium would yield partial derivatives of the energy density with respect to particle fractions, which are zero in beta equilibrium. Thus the energy density is in phase with the baryon density for small amplitude oscillations around beta equilibrium and it drops out of Eq. 10..
The energy lost by the oscillation is gained by the fluid element, and thus Eqs. 7 and 10 sum to zero and we find
@@
$P=P_{0}+\dfrac{\partial P}{\partial n_{B}}\bigg{|}_{T,x_{p},x_{\mu},x_{\pi}}\Re{(\delta n_{B}e^{i\omega t})}+\dfrac{\partial P}{\partial x_{p}}\bigg{|}_{T,n_{B},x_{\mu},x_{\pi}}\Re{(\delta x_{p}e^{i\omega t})}+\dfrac{\partial P}{\partial x_{\mu}}\bigg{|}_{T,n_{B},x_{p},x_{\pi}}\Re{(\delta x_{\mu}e^{i\omega t})}+\dfrac{\partial P}{\partial x_{\pi}}\bigg{|}_{T,n_{B},x_{p},x_{\mu}}\Re{(\delta x_{\pi}e^{i\omega t})}.$ (13)
Evidently (from Eqs. 4, 12, and 13)
$\Im{(\delta P)}=\dfrac{\partial P}{\partial x_{p}}\bigg{|}_{T,n_{B},x_{\mu},x_{\pi}}\Im{(\delta x_{p})}+\dfrac{\partial P}{\partial x_{\mu}}\bigg{|}_{T,n_{B},x_{p},x_{\pi}}\Im{(\delta x_{\mu})}+\dfrac{\partial P}{\partial x_{\pi}}\bigg{|}_{T,n_{B},x_{p},x_{\mu}}\Im{(\delta x_{\pi})},$ (14)
@@ III.1 Classification of chemical reactions in neutrino-trapped $npe\mu\pi^{-}$ matter
There are six classes [^2] Reactions within each class are related by “crossing” (e.g. direct Urca neutron decay and electron capture) or by the addition of a spectator particle to both sides of the reaction (e.g. direct Urca neutron decay and modified Urca neutron decay). of weak reactions that can occur in neutrino-trapped $npe\mu\pi$ matter, and therefore six chemical potential differences $\delta\mu_{i}$ that are zero in chemical equilibrium. However, three of them are redundant and can be written in terms of three independent equilibrating chemical potentials
@@
We assume in this analysis that particles are chemically equilibrated with their respective antiparticles, $\mu_{X}=-\mu_{\bar{X}}$. This is a good assumption for the hadrons and charged leptons, though for neutrinos it is merely a simplifying assumption [^3] If neutrinos and antineutrinos were not equilibrated with each other, as is likely the case except at quite large temperatures, we would have to assign distinct chemical potentials $\mu_{\nu_{e}},\mu_{\bar{\nu}_{e}},\mu_{\nu_{\mu}},\mu_{\bar{\nu}_{\mu}}$ and consider reactions such as $n+n\leftrightarrow n+n+\nu_{e}+\bar{\nu}_{e}$ or $\nu_{e}+\bar{\nu}_{e}\leftrightarrow\nu_{\mu}+\bar{\nu}_{\mu}$.. Given this, the six classes of reactions are listed in Table 1.
@@
6) $\delta\mu_{6}\equiv\mu_{\mu}+\mu_{\nu_{e}}-\mu_{e}-\mu_{\nu_{\mu}}=\delta\mu_{1}-\delta\mu_{2}$ $\mu^{-}\leftrightarrow e^{-}+\bar{\nu}_{e}+\nu_{\mu}$ $\lambda_{i}(\delta\mu_{1}-\delta\mu_{2})$
@@
The process $n\leftrightarrow p+\pi^{-}$ needs a spectator to proceed, but even the spectator reaction $n+n\leftrightarrow n+p+\pi^{-}$ is very fast compared to all other timescales in the problem as long as $T\gtrsim 1\text{ MeV}$, so in our numerical calculations we will take the $\lambda_{3}\rightarrow\infty$ limit and never have to calculate the rate of $n+n\leftrightarrow n+p+\pi^{-}$. The rate of this process, however, can be calculated using models such as the one developed in [57].